A market consists of three customers, A,B and C, whose individual demand equations are as follows:

                A:  P = 35 – 0.5QA

                B:  P = 50 – 0.25QB

                C:  P = 40 – 2.00Qc

The industry supply equation is given by Qs = 40 + 3.5P

a). Determine the equilibrium price and quantity.

b). Determine the amount that will be purchased by each individual


1
Expert's answer
2021-04-29T10:27:14-0400

1) Industry demand is the sum of demands of A, B, C. Firstly express demands as:

A: Qa=70-2P

B: Qb=200-4P

C: Qc=20-0.5P.

Then industry demand is Q=70-2P+200-4P+20-0.5P=290-6.5P.

Equilibrium price can be found from the equation

40+3.5P=290-6.5P;

P=25.

Equilibrium quantity is 40+3.5×25=127.5

b)Qa=70-50=20

Qb=200-100=100

Qc=20-12.5=7.5